OpenStudy (anonymous):

r=2/3t +y solve for t

OpenStudy (omnomnom):

\[r = 2\div 3t + y\]

OpenStudy (anonymous):

r -y = 2/3t t = (r-y)/(2/3) 

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

shall i do it step by step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\large \begin{align} r&=\frac{2}{3t}+y\\ r-y&=\frac{2}{3t}\\ 3t&=\frac{2}{r-y}\\ t&=\frac{2}{3(r-y)}\end{align}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

YOU ALREADY GOT HELP!!

OpenStudy (omnomnom):

xD lol nice